OpenStudy (anonymous):

Find the slope of the line passing through the points (2, 5) and (0, 4). (Points : 1) Option A: 9/2 Option B: 2/9 Option C: -9/2 Option D: -2/9

OpenStudy (anonymous):

HELPPPPP

OpenStudy (anonymous):

did you type the question or answers right? I'm not getting anything

OpenStudy (anonymous):

4-5/0-2 = 1/2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the 4 is supposed to be negative in the question

OpenStudy (anonymous):

option A 9/2

OpenStudy (anonymous):

-4-5/0-2 = -9/-2 = 9/2

OpenStudy (anonymous):

Thanks! ^_^